# Entanglement-swapping in generalised probabilistic theories, and iterated CHSH games

### v3

@@ Abstract
While there exist theories that have states “more strongly entangled” than quantum theory, in the sense that they show CHSH values above Tsirelson’s bound, all known examples of such theories have a strictly smaller set of measurements. Therefore, in tasks which require both bipartite states and measurements, they do not perform better than QM. One of the simplest information processing tasks involving both bipartite states and measurements is that of entanglement swapping. In this paper, we study entanglement swapping in generalised probabilistic theories (GPTs). In particular, we introduce the iterated CHSH game, which measures the power of a GPT to preserve non-classical correlations, in terms of the largest CHSH value obtainable after $n$ rounds of entanglement swapping. Our main result is the construction of a GPT that achieves a CHSH value of $4$ after an arbitrary number of rounds. This addresses a question about the optimality of quantum theory for such games recently raised in the literature. One challenge faced when treating this problem is that there seems to be no general framework for constructing GPTs in which entanglement swapping is a well-defined operation. Therefore, we introduce an algorithmic construction that turns a bipartite GPT into a multipartite GPT that supports entanglement swapping, if consistently possible.
@@ I Introduction
Therefore, by adding a step of entanglement swapping before playing the CHSH game, the set of possible correlations achievable in theories with such state spaces shrink. Quantum theory seems to strike the optimum between these competing notions, in that the CHSH violation of $2\sqrt{2}$ is preserved under entanglement swapping. And indeed, Weilenmann and Colbeck show that, for the adaptive CHSH game [22], the winning probability of any theory whose unipartite state and effect spaces are characterised by regular polygons [23] is upper bounded by the winning probability of quantum theory.
@@ II.1 Consistent Bipartite Theory
  • a convex cone $P\in V\otimes V$,
  • a convex cone $D\in(V\otimes V)^{*}$,
@@
Algorithm 1 Function: check consistency
@@
In the presentation of Algorithm 1, we have not specified in which format inputs like $V$ or $P$ are to be supplied, or how the checks should be performed. In this sense, it is a “template” for a concrete algorithm that depends on the mathematical properties of the input data. For example, if $P,D$ are polyhedral cones in $\mathbb{R}^{d}$, then $V$ can be represented by the integer $d$ and the cones by the facet inequalities. In this case, all tests in Algorithm 1 are linear programs. But more general situations also make sense, e.g. cones defined in terms of semi-definite constraints.
Algorithm 2 Induced theory
1:input: $(V,\underline{1},P,D,n\in\mathbb{N})$
2:if check consistency($V,\underline{1},P,D$) = “Inconsistent” then
3:     return “Inconsistent”
4:else
5:     $\mathcal{D}^{(1)}\leftarrow\mathrm{cone}\langle D,\underline{1}\rangle$
7:     $\mathcal{P}^{(2)}\leftarrow P$
8:     $\mathcal{D}^{(2)}\leftarrow D$
9:     if $n$ is even then
12:     else
15:     end if
16:     return $\mathcal{P}^{(n)},\mathcal{D}^{(n)}$
17:end if
Here $\dot{\otimes}$ stands for the minimal tensor product, defined as the conal hull over the tensor product of the respective sets:
$\mathcal{P}^{(n)}\dot{\otimes}\mathcal{P}^{(m)}:=\mathrm{cone}\big{(}\mathcal{P}^{(n)}\otimes\mathcal{P}^{(m)}\big{)}$ (15)
@@
If Algorithm 2 returns “inconsistent”, then one of the axioms in Def. 1 is violated already at the level of bipartite objects and so there is no consistent extension of the input.
We will now verify that if Algorithm 2 returns “consistent”, then there exists a theory as advertised.
@@ V.1 Setup
$\begin{gathered}|x_{\pm}\rangle\langle x_{\pm}|:=\tfrac{1}{2}(\mathbb{1}\pm\sigma_{1}),\\ |y_{\pm}\rangle\langle y_{\pm}|:=\tfrac{1}{2}(\mathbb{1}\pm\sigma_{2}),\\ |z_{\pm}\rangle\langle z_{\pm}|:=\tfrac{1}{2}(\mathbb{1}\pm\sigma_{3}).\end{gathered}$ (27)
Now perform a “uniform stretch in the equatorial plane of the Bloch sphere”, i.e., for some $r>1$, set
@@ V.3 Bipartite Theory
The data specified in Def. 5, can be consistently extended to a theory using Alg. 2.
@@
Therefore, none of the conditions required to detect inconsistency in Algorithm 1 are met. And hence, Algorithm 2 does not return “inconsistent”. ∎
@@ V.4 Iterated CHSH game
Given these choices one can verify (refer to Appendix. VIII.2) that the four output states obtained after each consecutive entanglement swap are the same (tabulated below). Each state is also equiprobable. The output state after $n$ rounds depends solely on the multiplicity of each of the four outcomes of $\mathcal{M}_{B_{i}}$, in the vector $\vec{b}=(b_{1},\cdots,b_{n})\in(\mathbb{Z}_{4})^{n}$. Therefore to obtain the proper correction one can convert the outcomes to binary and perform a bit-wise XOR (equivalent to finding the resultant element of the Klein four-group). The output state and correction corresponding to each $\mu\in\mathbb{Z}_{4}$ is as follows:
@@ VI Conclusion
As an outlook to future work, Ref. [20] suggests that theories should satisfy stronger symmetry conditions, which our construction does not have. In particular, our construction breaks the symmetry between $z$-observables and those on the equatorial plane. This raises two complementary questions for further work: (1) Are there natural, stronger conditions on multipartite correlations for which QM is indeed optimal? (2) Can one find a theory that beats QM in the iterated CHSH game and that is isotropic in the sense of having a transitive symmetry group action on all extremal effects?

### v4

@@ Abstract
While there exist theories that have states “more strongly entangled” than quantum theory, in the sense that they show CHSH values above Tsirelson’s bound, all known examples of such theories have a strictly smaller set of measurements. Therefore, in tasks which require both bipartite states and measurements, they do not perform better than QM. One of the simplest information processing tasks involving both bipartite states and measurements is that of entanglement swapping. In this paper, we study entanglement swapping in generalised probabilistic theories (GPTs). In particular, we introduce the iterated CHSH game, which measures the power of a GPT to preserve non-classical correlations, in terms of the largest CHSH value obtainable after $n$ rounds of entanglement swapping. Our main result is the construction of a GPT that achieves a CHSH value of $4$ after an arbitrary number of rounds. This addresses a question about the optimality of quantum theory for such games recently raised by Weilenmann and Colbeck. One challenge faced when treating this problem is that there seems to be no general framework for constructing GPTs in which entanglement swapping is a well-defined operation. Therefore, we introduce an algorithmic construction that turns a bipartite GPT into a multipartite GPT that supports entanglement swapping, if consistently possible.
@@ I Introduction
Therefore, by adding a step of entanglement swapping before playing the CHSH game, the set of possible correlations achievable in theories with such state spaces shrink (see [22, 23]). Quantum theory seems to strike the optimum between these competing notions, in that the CHSH violation of $2\sqrt{2}$ is preserved under entanglement swapping. And indeed, Weilenmann and Colbeck show that, for the adaptive CHSH game [24], the winning probability of any theory whose unipartite state and effect spaces are characterised by regular polygons [25] is upper bounded by the winning probability of quantum theory.
@@ II.1 Consistent Bipartite Theory
  • a convex cone $P\subset V\otimes V$,
  • a convex cone $D\subset(V\otimes V)^{*}$,
@@
Algorithm 1 Induced theory
1:input: $(V,\underline{1},P,D,n\in\mathbb{N})$
2:if check consistency($V,\underline{1},P,D$) = “Inconsistent” then
3:     return “Inconsistent”
4:else
5:     $\mathcal{D}^{(1)}\leftarrow\mathrm{cone}\langle D,\underline{1}\rangle$ partial contractions
7:     $\mathcal{P}^{(2)}\leftarrow P$
8:     $\mathcal{D}^{(2)}\leftarrow D$
9:     if $n$ is even then
12:     else
15:     end if
16:     return $\mathcal{P}^{(n)},\mathcal{D}^{(n)}$
17:end if
@@
In the presentation of Algorithm 1, $\dot{\otimes}$ stands for the minimal tensor product, defined as the conal hull over the tensor product of the respective sets:
$\mathcal{P}^{(n)}\dot{\otimes}\mathcal{P}^{(m)}:=\mathrm{cone}\big{(}\mathcal{P}^{(n)}\otimes\mathcal{P}^{(m)}\big{)}.$ (15)
Additionally, we have not specified in which format inputs like $V$ or $P$ are to be supplied, or how the checks should be performed. In this sense, it is a “template” for a concrete algorithm that depends on the mathematical properties of the input data. For example, if $P,D$ are polyhedral cones in $\mathbb{R}^{d}$, then $V$ can be represented by the integer $d$ and the cones by the facet inequalities. In this case, all tests in Algorithm 1 are linear programs. But more general situations also make sense, e.g. cones defined in terms of semi-definite constraints.
@@
If the function check consistency in Algorithm 1 returns “inconsistent”, then one of the axioms in Def. 1 is violated already at the level of bipartite objects and so there is no consistent extension of the input.
We will now verify that if the function check consistency in Algorithm 1 returns “consistent”, then there exists a theory as advertised.
@@ V.1 Setup
$\begin{gathered}|x_{\pm}\rangle\langle x_{\pm}|:=\tfrac{1}{2}(\mathbb{1}\pm\sigma_{1}),\\ |y_{\pm}\rangle\langle y_{\pm}|:=\tfrac{1}{2}(\mathbb{1}\pm\sigma_{2}),\\ |z_{\pm}\rangle\langle z_{\pm}|:=\tfrac{1}{2}(\mathbb{1}\pm\sigma_{3}),\end{gathered}$ (27)
where $\sigma_{1},\sigma_{2},\sigma_{3}$ are the Pauli matrices. Now perform a “uniform stretch in the equatorial plane of the Bloch sphere”, i.e., for some $r>1$, set
@@ V.3 Bipartite Theory
The data specified in Def. 5, can be consistently extended to a theory using Alg. 1.
@@
Therefore, none of the conditions required to detect inconsistency in Algorithm 1 are met. And hence, it does not return “inconsistent”. ∎
@@ V.4 Iterated CHSH game
Remark 7.
It is worth pointing out here that the choice of correlators being only on the $xy$-plane is no coincidence. For any situation with a $z$-measurement, oblate stabilizer theory no longer has such a strong CHSH violation.
@@
Given these choices one can verify (refer to Appendix. VIII.2) that the four output states obtained after each consecutive entanglement swap are the same. Each state is also equiprobable. The output state after $n$ rounds depends solely on the multiplicity of each of the four outcomes of $\mathcal{M}_{B_{i}}$, in the vector $\vec{b}=(b_{1},\cdots,b_{n})\in(\mathbb{Z}_{4})^{n}$. Therefore to obtain the proper correction one can convert the outcomes to binary and perform a bit-wise XOR (equivalent to finding the resultant element of the Klein four-group). The output state and correction corresponding to each $\mu\in\mathbb{Z}_{4}$ is tabulated in Tab. 1.
@@
Table 1: The output state and correction corresponding to each outcome of Bob’s measurement.
@@ VI Conclusion and Outlook
As an outlook to future work, Ref. [20] suggests that theories should satisfy stronger symmetry conditions, i.e., “…for any state and set of local measurements, if the local outcome probabilities are permuted, then there is a state that achieves these permuted correlations under the same measurements”. If this is interpreted as an invarience under permutation of sub-systems, then oblate stabilizer theory satisfies this requirement. If instead we interpret this as a symmetry under permutation of extremal effects, then our construction fails to satisfy this requirement. In particular, our construction breaks the symmetry between $z$-observables and those on the equatorial plane. This raises two complementary questions for further work: (1) Are there natural, stronger conditions on multipartite correlations for which QM is indeed optimal? (2) Can one find a theory that beats QM in the iterated CHSH game and that is isotropic in the sense of having a transitive symmetry group action on all extremal effects?